Consider the following information for Evenflow Power Co., Debt: 6,000 5.5 percent coupon bonds outstanding, $1,000 par value, 22 years to maturity, selling for 102 percent of par; the bonds make semiannual payments. Common stock: 144,000 shares outstanding, selling for $60 per share; the beta is 1.06. Preferred stock: 21,500 shares of 4.5 percent preferred stock outstanding, currently selling for $104 per share. Market: 7 percent market risk premium and 3.5 percent risk-free rate. Assume the company's tax rate is 32 percent. Required: Find the WACC. (Do not round your intermediate calculations.)
Answer #1

Value of debt = 6000*1020 = 6120000

Value of common equity = 144000*60 = 8640000

Value of preferred stock = 21500*104 = 2236000

Total value = 6120000 + 8640000 + 2236000 = 16996000

Cost of debt:

PV = 1020

N = 22*2 = 44

PMT = 0.055*1000/2 = 27.5

FV = 1000

R = Rate(N,PMT,PV,FV)
Rate(44,-27.5,1020,-1000) = 2.67%

Annual cost of debt = $2.67%*2 = 5.34%

Cost of equity:

Cost of equity = Rf + beta*risk premium = 3.5%+1.06*7% = 10.92%

Cost of preferred stock = Annual dividend/price = 4.5/104 = 4.33%

WACC = rD (1- Tc )*( D / V )+ rE *( E / V )

Where...

rD = The required return of the firm's Debt financing
(1-Tc) = The Tax adjustment for interest expense
(D/V) = (Debt/Total Value)
rE= the firm's cost of equity
(E/V) = (Equity/Total Value)

WACC = 6120000/16996000 * 0.0534 * (1-0.32) + 8640000/16996000 * 0.1092 + 2236000/16996000 * 0.0433 = 0.0743 = 7.43%